OpenStudy (anonymous):

Max is 5 years older than Paulette. Next year he will be twice as old as she will be. How old is each now?

OpenStudy (anonymous):

let paulette age=x max=x+5

OpenStudy (anonymous):

now after 1 year 2(x+1)=x+6

OpenStudy (anonymous):

2x+2=x+6 x=4

OpenStudy (anonymous):

hence paulette age=4 years and max's age =9years

OpenStudy (anonymous):

:)

Directrix (directrix):

Let P represent Paulette's age now Let M represent Max's age now. Max is 5 years older than Paulette --> so P + 5 = M Max's Age next year: M + 1 Paulette's age next year: P + 1 Next year he will be twice as old as she will be. M + 1 = 2 ( P + 1) but M = P + 5 so P + 5 + 1 = 2 ( P + 1) P + 6 = 2P + 2 P = 4 --> Paulette's age now So, what is Max's age now? M = P + 5. @velasoairajoy

OpenStudy (anonymous):

M = P + 5 + 5 = 9p+1) 1 = p - 3 p= 4 m = 4 + 5 = 9 Max is 9 yrs-old. Paullete is 4 yrs-old